Check my work Be sure to answer all parts. A chemist mixes solid AgCL CuCh, and Mga2 ǐn enough water to give a final volume of 50.0 mL (o) With ions shown as spheres and solvent molecules omitted for clarity, which of the following best represents the resulting mixture? O Ag Cl Cu O Mg K Prev 7 of 30 Next>
media%2F41c%2F41cb5c91-5c45-4184-b0e2-5b
Answer #1

a) Screen A represents the correct resulting mixture.

This is because AgCl forms precipitate immediately on mixing with water where as the other ions will be in dissolved state.

Screen -B is wrong because it has one excess chloride ion (odd numbered) which donot balance the charges of cations.

Screen c and D are wrong because Magnesium salts are readily soluble in water and donot form precipitate.

b)

In screen A we can count the free spheres (which did not form precipitate) as 2 red, 2 purple and 8 green. total 12 spheres.

so total ionic concentration = 12 x 2.0 x 10-3 mol = 0.024 mol

volume of water taken = 50.0 ml = 0.05 L

Molarity = no. of moles / Volume in L = 0.024 / 0.05 = 0.48 M

C) Total mass of solid precipitate can be calculated as follows

white spheres 4 and green spheres 4 total 8 spheres precipitated.

so concentration = 8 x 2.0 x 10-3 = 0.016 mol

Molar mass of AgCl = 143.32 g/mol

so 1 mole has mass of 143.32 g

0.016 mole has mass = 143.32 x 0.016 =2.29312 g = 2.29 g